OpenStudy (anonymous):

WILL GIVE MEDAL Choose the correct slope of the line that passes through the points (-4, 8) and (-3, -6).

OpenStudy (anonymous):

I also have this question Given the equation y - 3 = 1/2 (x + 6) in point-slope form, identify the equation of the same line in standard form.

OpenStudy (anonymous):

1. Is -14/3

OpenStudy (anonymous):

the choices are 7 -14 0 14

OpenStudy (anonymous):

@Niaa08

OpenStudy (anonymous):

I'm Here Hold On

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

The Answer Is -14

OpenStudy (anonymous):

thanks wb the second question

OpenStudy (anonymous):

I'm doing it now

OpenStudy (anonymous):

What are the answer choices

OpenStudy (anonymous):

x - 2y + -12 y = 1/2x y = 1/2x + 6 y = 1/2x + 9

OpenStudy (anonymous):

The answer is 1x-2y=12

OpenStudy (anonymous):

huh that's not any of the choices.

OpenStudy (anonymous):

Sorry I Have A Habit Of Putting A 1 in Front Of The X's But It's The Same Thing X-2y=12

OpenStudy (anonymous):

That's still not one of the answers.

OpenStudy (anonymous):

Your Question said in standard form X+y= Is Standard form The choices you gave me were in slope intercept

OpenStudy (anonymous):

OH WAIT never mind. Thank you @Niaa08

OpenStudy (anonymous):

Your Welcome Do You Need Any more help